OpenStudy (mven):

The radius of one sphere is twice as great as the radius of a second sphere. Find the ratio of their volumes.

OpenStudy (mven):

@iGreen Help please!!

OpenStudy (anonymous):

What's the equation to calculate the volume of a sphere if you know the radius?

OpenStudy (mven):

4pir^2

OpenStudy (mven):

I got 8 but that was wrong

OpenStudy (anonymous):

Well, I guess if you use any arbitrary number for the radius, and calculate the volume, and then do the same for a radius twice as big, and then divide the volumes from each other, that would do the trick?

OpenStudy (mven):

got 1:2 ?

OpenStudy (anonymous):

8 is pretty close actually

OpenStudy (mven):

it said that it was wrong. I thaught 8 too. otherwise it gives me 1/2 1/4 1/8 1/16 I dont think that 1/8 and 1/4 would work.

OpenStudy (mven):

it said 1/8 was wrong anyway

OpenStudy (anonymous):

Ah right, 1/8 as it's a ratio.

OpenStudy (anonymous):

So the one is 8 times bigger to the other, as a ratio you write 1/8 (as in, if the one has a volume of 1, the other has a volume of 8)

OpenStudy (mven):

ah, Makes sense. so one being 2 times larger as the other the ratio would be 1/2 ?

OpenStudy (anonymous):

True, but note that if the radius has a ratio of 1/2 the volume has a ratio of 1/8 because \[\frac{ 4 }{ 3 }*\pi* r^3 = x*\frac{ 4 }{ 3 }*\pi*(2r)^3 \] If you solve this either r=0 or x=1/8

OpenStudy (anonymous):

That is because the radius is to the power of 3 and 2^3 is 8

OpenStudy (mven):

ahh, Makes sense. Thanks for the help!!